OpenStudy (anonymous):

steps on how to evaluate this limit lim as x approaches 0 (e^x) -1/ sin(13x)

OpenStudy (tkhunny):

Step #1 is to write clearly. Do you REALLY mean what you have written? \((e^{x}) - \dfrac{1}{\sin(13x)}\) Or did you mean \(\dfrac{e^{x} - 1}{\sin(13x)}\)?

OpenStudy (anonymous):

oh, i'm sorry i should i have written it more clearly. the bottom one! sorry about that..

OpenStudy (anonymous):

have you tried l'hospital's rule?

OpenStudy (anonymous):

that's the part i don't understand how to do.. :(

OpenStudy (anonymous):

awh, ok do you know how to take the derivative?

OpenStudy (tkhunny):

First, make sure we have an Indeterminate Form. Don't just use L'Hospital's rule willy-nilly.

OpenStudy (tkhunny):

I'm tempted just to rewrite it. \(\dfrac{e^{x}-1}{\sin(13x)} =\dfrac{13x}{\sin(13x)}\cdot\dfrac{e^{x} - 1}{13x}\) The expression on the right is quite a bit easier to fathom.

OpenStudy (anonymous):

would i take the derivative of the expression on the right?

OpenStudy (anonymous):

I'd just take the derivative of what they gave you; e^x - 1 and sin(13x)

OpenStudy (anonymous):

d (e^x - 1)dx = ?

OpenStudy (anonymous):

would it be e^x ?

OpenStudy (anonymous):

for sin(13x) would the derivative be cos(13x) ?

OpenStudy (anonymous):

e^x is correct for sin(13x) you'll need to use 'u-substitution' it would look like this: d(sin(13x))dx, let u = 13x, then du = 13, and d(sin(u)) = cos(u)du cos(u)du = cos(13x)*13 or 13cos(13x)

OpenStudy (anonymous):

now take the limit of this: \[\lim_{x \rightarrow 0}\frac{ e^x }{ 13\cos(13x) }\]

OpenStudy (anonymous):

would it be 1/13 ?

OpenStudy (tkhunny):

Under what circumstances would it be something else?

OpenStudy (anonymous):

haha i just wanted to make sure... thank you everyone!